The equation of two variables x and y are follows: 3x+2y-26=0;6x+y-31=0
Find a)mean
b) regression correlation
1
Expert's answer
2019-05-27T13:13:16-0400
"3x+2y-26=0""6x+y-31=0"

Solving these equations simultaneously, we get 


"6x+4y-52=0""6x+y-31=0"

"3x+2y-26=0""3y-21=0"

"3x=-2(7)+26""y=7"

Then "\overline{x}=4, \overline{y}=7."

Let "3x+2y-26=0" be the regression line of X on Y and the other line as Y on X. Then


"x=-{2 \over 3}y+{26 \over 3}=>b_{xy}=-{2 \over 3}"

"y=-6x+31=>b_{yx}=-6"

But "r^2=b_{xy}\cdot b_{yx}=4" which cannot be true, because "-1\leq r\leq 1."

So we change our assumptions i.e., the line "6x+y-31=0" be the regression line of X on Y and the other line as Y on X. Then


"x=-{1 \over 6}y+{31 \over 6}=>b_{xy}=-{1 \over 6}"

"y=-{3 \over 2}x+13=>b_{yx}=-{3 \over 2}"

"r^2=b_{xy}\cdot b_{yx}=(-{1 \over 6})(-{3 \over 2})={1 \over 4}"

As the regression coefficients have negative signs, we have to consider negative value of "r."


"r=-\sqrt{{1 \over 4}}=-{1 \over 2}"

"a)\ \ \ \overline{x}=4, \overline{y}=7""b)\ \ \ r=-0.5 \ \ \ \ \"